OpenStudy (anonymous):

How does the zero product property help you find the x-intercept of y = x² - x - 12?

OpenStudy (anonymous):

To find x intercept plug y = 0 into given equation So we get \[x^2-x-12=0\] Now use quadratic formula or AC method

OpenStudy (anonymous):

Because if we can factor it into something like: \[ (x-a)(x-b) =0 \]Then if \(x-a=0\) the whole thing is zero. The same goes for \(x-b=0\).

OpenStudy (anonymous):

Yes but before this we have to find out groups .....